Can someone please help me with this problem: x^2+5x-6=0 I need to factor completely and I am lost. Thank you in advance!
Answer by stanbon(48569)

x^2+5x-6=0
(x+6)(x-1)=0
x=-6 or x=1
Cheers,
Stan H.